OpenStudy (anonymous):

how do I convert 32.5% to a decimal & a fraction?

OpenStudy (anonymous):

just divide number with 100. 32.5/100. and you will get your fraction value

OpenStudy (anonymous):

sweet! Thanks guys!!

OpenStudy (anonymous):

Any idea how to convert a fraction with a whole number, like 1 3/5

OpenStudy (anonymous):

i will show you this from a simple example . like i have 0.25, so to get whole number. i will do 25/100 = 1/4. because 100 can be divide by 25 completely . numbers are not always completely divided.like if we have 40/100 = firstly 0 will cut by 0. then we 4/10, we can divide it on 2. after divide we will 2/5.

OpenStudy (anonymous):

Im confused lol :(

OpenStudy (anonymous):

Ihave to convert 1 3/5 to a decimal & a fraction how do u convert to a decimal?

OpenStudy (anonymous):

... when using a whold number

OpenStudy (anonymous):



OpenStudy (anonymous):

first you multiply 5 with 1 and then add 3 in it. so , it will be (5*1+3)/5 = 8/5.

OpenStudy (anonymous):

OMG!!ur awesome!!thanks sooooo much! I remember now!! lol

OpenStudy (anonymous):

its, you get now how to solve it. practice it as much as you can , it will be good for you in future.@WhatTheWhat

OpenStudy (anonymous):

will do for sure!! Thanks again!

OpenStudy (rina.r):